# HSIGene: A Foundation Model For Hyperspectral Image Generation

### v1

@@ III-A1 HSI data collection
III-A2 Condition generation
In our work, we consider six control conditions for HSI synthesis, including holistically-nested edge detection (HED), segmentation, sketch, multiscale line segment detection (MLSD), content and text. We use the RGB bands of HSIs to generate conditions since most existing models are designed for RGB images. A more detailed description of the conditions is provided in the following.
@@ III-A2 Condition generation
Sketch: The sketch drawings of HSI data is obtained using the model proposed in [52], which offer a simplified representation that reduces an image to its essential lines and contours defining the object shapes within the image.
MLSD (Multiscale Line Segment Detection): We use the model proposed in [53] to achieve line segment detection of HSI data, providing the straight paths between different endpoints.
@@
Text: To enable text-to-image generation, we annotate 1k training images with labels corresponding to the categories of the images. The labeling process involved assigning each image to one of four specific categories: farmland, city building, architecture, and wasteland. The labeled text enables the model to support category-guided synthesis.
III-B Controllable Generation Model
@@
where $z_{t}$ represents the noisy data at time step $t$ and $\alpha_{t}$ is a variance schedule parameter that controls the amount of noise added at each step. In the reverse process, with the low-resolution image as guidance, the model starts from the noise distribution, gradually removes noise to generate high-resolution samples. In our work, the sampling method proposed in Denoising Diffusion Implicit Models (DDIM) [25] is adopted since the method allows for non-sequential steps in the sampling process, leading to significant speedups in generating samples. The reverse process is given by:
@@
where $\epsilon_{t}$ is the Gaussian noise added to the latent representation at time step $t$, $\hat{\epsilon}_{t}$ is the noise predicted by the model, and $z_{0}$ is the latent representation of high-quality image. After training, the model can transform random noise into realistic high-resolution samples by sequentially denoising at each step with multiple conditions such as sketch and segmentation.
The conditions are fed into the ControlNet encoder to ensure that the generated content aligns with the original low-resolution image. Specifically, the conditions are first fed into the feature extractor module, which consists of a series of convolutional layers to capture spatial hierarchies and semantic features. Then, the outputs from the zero-convolution modules are concatenated with those from the UNet in the diffusion model to effectively integrate the conditions’ information, ensuring that the generated images reflect the multi-conditional information.
III-C HSI Data Augmentation
III-C1 Overall framework
To further improve the model’s ability to handle various conditions, we propose to augment the training data with additional samples. The augmented data are expected to satisfy two constraints: Firstly, the augmented data should exhibit desirable perceptual quality. Secondly, the spectral distribution of the augmented data should closely match that of real HSIs. Therefore, we propose to perform spatial super-resolution for real HSIs with diffusion models as shown in Fig. 1(b). By doing so, diverse training samples can be extracted by cropping the high-resolution HSIs. The augmented samples adhere to the two principles mentioned above as the super-resolution operation preserves spectral distribution and the diffusion models could ensure the high perceptual quality of the super-resolved results. To further improve the perceptual quality of the augmented data, we propose a novel two-stage framework for HSI super-resolution as shown in Fig. 4, which takes advantage of abundant high-resolution remote sensing images to enhance super-resolution performance. In this framework, the RGB bands of HSIs are super-resolved first, and then the RGB details are utilized as a priori information to enhance the HSI super-resolution performance. Specifically, we first collect abundant high-resolution remote sensing images whose content is highly similar to real HSIs using Google Earth Engine [55] and a diffusion-based RGB super-resolution network (DSRNet) is trained. This network then enhances the resolution of the RGB bands of HSIs, producing high-resolution RGB bands. Finally, by using enhanced RGB bands as guidance, a novel Rectangular Guided Attention Network (RGAN) is proposed to generate high-resolution HSIs. More details of the networks are introduced in the following.
@@ III-C1 Overall framework
III-C2 Diffusion based super-resolution
As diffusion models have achieved great success in the field of image generation, an increasing number of studies have applied diffusion models to image restoration tasks to enhance the perceptual quality of the restored results [56, 57, 58]. In our work, we employ the diffusion-based super-resolution network (DSRNet) which is highly based on the framework introduced in Sec. III-B for RGB super-resolution. Specifically, we added an upsampling module before the feature extractor and regarded low-resolution images as conditions. The model is employed to obtain high-resolution RGB bands of HSIs, which provide image texture details for HSI super-resolution.
@@ III-C2 Diffusion based super-resolution
III-C3 RGB-modality guided super-resolution
The RGB-guided attention super-resolution network, i.e. RGAN, is designed to obtain high-resolution HSIs with high-resolution RGB-modality as auxiliary information. As demonstrated in Fig. 5, RGAN is composed of multiple guided attention layers (GALs). Each GAL consists of self attention layer (SAL), cross attention layer (CAL), spectral attention layer (SpecAL) and feed forward layer (FFD). We adopt the rectangle self-attention (RA) module and spectral enhancement (SE) module proposed in [59] as the main components of CAL and SpecAL to extract spatial and spectral correlations efficiently. FFD is composed of a normalization layer and two linear layers with a ReLU activation. In addition, inspired by the rectangle self-attention proposed in [59], a novel rectangular cross-attention (RCA) module is designed to aggregate the mutual contextual information between RGB and HSI efficiently.
The details of our proposed RCA module is illustrated in Fig. 6. The cross-attention is conducted in vertical and horizontal rectangles separately. Specifically, let $\boldsymbol{Z}_{1}\in\mathbb{R}^{H\times W\times C},\boldsymbol{Z}_{2}\in\mathbb{R}^{H\times W\times C}$ denote the input features, the query, key and value features of the two inputs are calculated as
@@ III-C3 RGB-modality guided super-resolution
$\small\begin{split}&\boldsymbol{Q}_{1}^{1},\boldsymbol{Q}_{1}^{2}=\text{Split}(\boldsymbol{Q}_{1}),\boldsymbol{K}_{1}^{1},\boldsymbol{K}_{1}^{2}=\text{Split}(\boldsymbol{K}_{1}),\boldsymbol{V}_{1}^{1},\boldsymbol{V}_{1}^{2}=\text{Split}(\boldsymbol{V}_{1})\\ &\boldsymbol{Q}_{2}^{1},\boldsymbol{Q}_{2}^{2}=\text{Split}(\boldsymbol{Q}_{2}),\boldsymbol{K}_{2}^{1},\boldsymbol{K}_{2}^{2}=\text{Split}(\boldsymbol{K}_{2}),\boldsymbol{V}_{2}^{1},\boldsymbol{V}_{2}^{2}=\text{Split}(\boldsymbol{V}_{2})\\ \end{split}$ (5)
Then, $\boldsymbol{Q}_{1}^{1}$,$\boldsymbol{Q}_{2}^{1}$,$\boldsymbol{K}_{1}^{1}$,$\boldsymbol{K}_{2}^{1}$,$\boldsymbol{V}_{1}^{1}$,$\boldsymbol{V}_{2}^{1}$ are partitioned into horizontal windows and $\boldsymbol{Q}_{1}^{2}$,$\boldsymbol{Q}_{2}^{2}$,$\boldsymbol{K}_{1}^{2}$,$\boldsymbol{K}_{2}^{2}$,$\boldsymbol{V}_{1}^{2}$,$\boldsymbol{V}_{2}^{2}$ are partitioned into vertical windows, respectively. As shown in Fig. 7, given the size of the rectangular window is $[h,w]$, the input features is partitioned into non-overlapping rectangular patches, and the size of each patch is $[h,w,\frac{C}{2}]$. In the case of horizontal partition, the width $w$ is greater than the height $h$ whereas in the case of vertical partition the height $h$ is greater than the width $w$. Next, these windows are fed into the multi-head cross-attention (MCA) module to integrate the features between the RGB modality and HSI modality. More precisely, the cross attention is calculated as
@@
Multi-head attention mechanism is also adopted in the RCA module, which indicates that we employ several groups of parameters to conduct the cross-attention, and the results are then combined to capture a richer representation of the data.
In summary, the input feature maps of HSI and RGB are divided into vertical and horizontal rectangular windows. Cross-attention is then applied between the vertical windows of the HSI and RGB feature maps, as well as between the horizontal windows. This rectangular attention mechanism enables the network to transfer the image details from RGB modality to HSI modality effectively and efficiently.
@@
Figure 8: Visualization results of synthesized HSIs which are generated under single condition. Our proposed HSIGene supports image generation under single condition including HED, segmentation, sketch, MLSD, content and text.
IV Experiments
IV-A Experimental Settings
IV-A1 Synthesis experiments setting
@@
The datasets used for HSI synthesis are listed in Table II and the details are illustrated in Sec. III-A. Considering the spectral discrepancy between different datasets, we align their wavelength to 400-1000nm with 48 bands utilizing linear interpolation which could ensure spectral information preservation [60]. The datasets are cropped to size $256\times 256$ with stride 128, resulting in approximately 7k training images. For data augmentation, we employ the method introduced in Sec. III to obtain high-resolution images of Heihe dataset with upscale factor ×$\times 2$ and Chikusei dataset with upscale factor ×$\times 4$. By cropping the super-resolved images of size $256\times 256$ with stride 128, we obtain 40k training images in total. To validate the robustness and generalization ability of our model, AID [61] dataset which exhibits scenes similar to real HSIs [27] is used for evaluation. The AID dataset is also cropped with a stride of 128, resulting in 1k image patches of size $256\times 256$. For each patch, we generate control conditions using the method introduced in Sec. III-A which are subsequently employed to synthesize HSIs. Finally, the quality of the generated images is evaluated to assess the model’s ability to generate realistic data.
@@
We use the AdamW optimizer [62] with a learning rate of $10^{-5}$ to train the HSI synthesis diffusion model for 100k iterations. The batch size is set as 16. There are 1.5 billion parameters in total for our model, ensuring desirable generation under scaling laws. In addition, we initialize the parameters of the UNet component from the pretrained weights provided by [36], ensuring that the diffusion model adapts effectively to HSI synthesis.
IV-A2 Super-resolution experiments setting
@@
Approximately 55k high-resolution remote sensing images obtained from Google Earth Engine [55] are used to train the DSRNet described in Sec.III-C2. We use real HSI data (i.e., the 7k cropped image patches) and select the RGB bands as auxiliary guidance to train the RGAN model proposed in III-C3. The training pairs are obtained by downsampling existing real HSIs with scales of 2 and 4 using the strategy proposed in [63].
@@
We adopt the same training strategy of the HSI synthesis model for the training of DSRNet. For the training of the RGAN model, the learning rate is set to $10^{-4}$ with AdamW optimizer and the learning rate finally decreases to $10^{-5}$ with a cosine annealing strategy [64]. The batch size is set as 2 and the total epoch number is set as 30. L1 loss is used to train the network. To evaluate the image quality of the augmented data, we randomly select 100 real images from Heihe dataset and compare the quality of these images super-resolved by different methods. Six super-resolution methods including MCNet [38], SSPSR [40], Bi-3DQRNN [39], SwinIR [65], ESSAformer [41], DSTrans [66] are taken for comparison. For fair comparison, all competitive models are trained in the same setting.
@@ IV-A3 Evaluation metrics
IV-B HSI synthesis results
IV-B1 Qualitative results
@@
In this section, we demonstrate the effectiveness of our model in achieving both single-condition and multi-condition control. The results of single condition generation are illustrated in Fig. 8. As shown in the figure, the generated images are consistent with the specified condition but also exhibit pixel-wise differences. For example, given the sketch of farmland, the generated images demonstrate highly similar structure while the image contents include buildings and wasteland. By employing conditions that indicate textural attributes, such as HED or sketch, the generated images successfully maintain the structure of the input conditions. On the other hand, when the control condition is related to the image content, the model generates images that align closely with the reference content and exhibit various structures. However, single condition control fails to guarantee simultaneous control over both the structure and content of the generated images. In addition, the model under single-condition constraints is unable to integrate multiple forms of guidance to achieve a more precise and coherent control.
@@
The results of multi-condition generation are illustrated in Fig. 9. Comparing with single-condition generation, more conditions enable more sophisticated and accurate generation. For example, as shown in Fig. 9, given conditions of HED and content indicating farmland, the generated images not only align with the structural prerequisites defined by the HED condition but also accurately reflect the content requirement. The multi-condition generation capability of our model offers significant advantages in producing images that are both structurally accurate and content-specific. Our model effectively handles different modalities and generates high-quality images, verifying the effectiveness and robustness of our model.
@@
In addition, the visualization comparison between the generated and real spectral profiles is provided in Fig. 10. We select some generated HSIs whose content is similar to real HSIs, and for these images, we plot the spectral profiles at the same point for analysis. It can be seen that the generated spectral profiles are consistent with real spectral profiles, proving the effectiveness and usability of our proposed model in HSI synthesis. We also provide t-SNE [78] visualizations of the spectra generated by UBF [27], UnmixDiff [28], and our proposed method as shown in Fig. 11. It can be seen that the distribution of spectra generated by our method covers the real distribution more effectively. UBF generates spectra with significant differences from the real spectra due to the discrepancy between RGB modality image content and HSI content. UnmixDiff, on the other hand, generates abundance maps rather than directly generating HSIs, leading to performance limitations imposed by the unmixing network. Furthermore, UnmixDiff has a smaller model size compared to our model, which restricts its generation capabilities. In contrast, the latent diffusion technique that we employ in this work, enables us to train a larger model, leading to better performance.

### v2

@@ 3.1.1 HSI data collection
3.1.2 Condition generation
In our work, we consider six control conditions for HSI synthesis, including holistically nested edge detection (HED), segmentation, sketch, multiscale line segment detection (MLSD), content and text. We use the RGB bands of HSIs to generate conditions since most existing models are designed for RGB images. A more detailed description of the conditions is provided in the following.
@@ 3.1.2 Condition generation
Sketch: The sketch drawings of HSI data are obtained using the model proposed in [40], which offers a simplified representation that reduces an image to its essential lines and contours defining the object shapes within the image.
MLSD (Multiscale Line Segment Detection): We use the model proposed in [41] to achieve line segment detection of HSI data, providing straight paths between different endpoints.
@@
Text: To enable text-to-image generation, we annotate 1k training images with labels corresponding to the categories of the images. The labelling process involved assigning each image to one of four specific categories: farmland, city building, architecture, and wasteland. The labelled text enables the model to support category-guided synthesis.
3.2 Controllable Generation Model
@@
where $z_{t}$ represents the noisy data at time step $t$ and $\alpha_{t}$ is a variance schedule parameter that controls the amount of noise added at each step. In the reverse process, with the low-resolution image as guidance, the model starts from the noise distribution and gradually removes noise to generate high-resolution samples. In our work, the sampling method proposed in Denoising Diffusion Implicit Models (DDIM) [17] is adopted since the method allows for non-sequential steps in the sampling process, leading to significant speedups in generating samples. The reverse process is given by:
@@
where $\epsilon_{t}$ is the Gaussian noise added to the latent representation at time step $t$, $\hat{\epsilon}_{t}$ is the noise predicted by the model, and $z_{0}$ is the latent representation of the high-quality image. After training, the model is able to transform random noise into realistic high-resolution samples by sequential denoising at each step with multiple conditions such as sketch and segmentation.
The conditions are fed into the ControlNet encoder to ensure that the generated content aligns with the original low-resolution image. Specifically, the conditions are first fed into the feature extractor module, which consists of a series of convolutional layers to capture spatial hierarchies and semantic features. Then, the outputs from the zero-convolution modules are concatenated with those from the UNet in the diffusion model to effectively integrate the conditions, ensuring that the generated images reflect the multi-conditional information.
3.3 HSI Data Augmentation
3.3.1 Overall framework
To further improve the model’s ability to handle various conditions, we propose to augment the training data with additional samples. The augmented data are expected to satisfy two constraints: Firstly, the augmented data should exhibit desirable perceptual quality. Secondly, the spectral distribution of the augmented data should closely match that of real HSIs. Therefore, we propose to perform spatial super-resolution for real HSIs with diffusion models as shown in Fig. 1(b). By doing so, diverse training samples can be extracted by cropping the high-resolution HSIs. The augmented samples adhere to the two principles mentioned above as the super-resolution operation preserves spectral distribution and the diffusion models could ensure the high perceptual quality of the super-resolved results. To further improve the perceptual quality of the augmented data, we propose a novel two-stage framework for HSI super-resolution as shown in Fig. 4, which takes advantage of abundant high-resolution remote sensing images to enhance super-resolution performance. In this framework, the RGB bands of HSIs are super-resolved first, and then the RGB details are utilized as a priori information to enhance the HSI super-resolution performance. Specifically, we first collect abundant high-resolution remote sensing images from Google Earth Engine [43] by matching their geographic coordinates with those of real HSIs, and thus the content of collected images is highly similar to real HSIs. Then, a diffusion-based RGB super-resolution network (DSRNet) is trained. This network enhances the resolution of the RGB bands of HSIs, producing high-resolution RGB bands. Finally, by using enhanced RGB bands as guidance, a novel Rectangular Guided Attention Network (RGAN) is proposed to generate high-resolution HSIs. More details of the networks are introduced in the following.
@@ 3.3.1 Overall framework
3.3.2 Diffusion based super-resolution
As diffusion models have achieved great success in the field of image generation, an increasing number of studies have applied diffusion models to image restoration tasks to enhance the perceptual quality of the restored results [44, 45, 46]. In our work, we employ the diffusion-based super-resolution network (DSRNet) which is highly based on the framework introduced in Sec. 3.2 for RGB super-resolution. Specifically, we added an upsampling module before the feature extractor and regarded low-resolution images as conditions. The model is employed to obtain high-resolution RGB bands of HSIs, which provide image texture details for HSI super-resolution.
@@ 3.3.2 Diffusion based super-resolution
3.3.3 RGB-modality guided super-resolution
The RGB-guided attention super-resolution network, i.e. RGAN, is designed to obtain high-resolution HSIs with high-resolution RGB-modality as auxiliary information. As demonstrated in Fig. 5, RGAN is composed of multiple guided attention layers (GALs). Each GAL consists of a self-attention layer (SAL), cross-attention layer (CAL), spectral attention layer (SpecAL) and feed-forward layer (FFD). We adopt the rectangle self-attention (RA) module and spectral enhancement (SE) module proposed in [47] as the main components of CAL and SpecAL to extract spatial and spectral correlations efficiently. FFD is composed of a normalization layer and two linear layers with a ReLU activation. In addition, inspired by the rectangle self-attention proposed in [47], a novel rectangular cross-attention (RCA) module is designed to aggregate the mutual contextual information between RGB and HSI efficiently.
The details of our proposed RCA module are illustrated in Fig. 6. The cross-attention is conducted in vertical and horizontal rectangles separately. Specifically, let $\boldsymbol{Z}_{1}\in\mathbb{R}^{H\times W\times C}$ and $\boldsymbol{Z}_{2}\in\mathbb{R}^{H\times W\times C}$ denote the input features, the query, key and value features of the two inputs can be calculated as
@@ 3.3.3 RGB-modality guided super-resolution
$\small\begin{split}&\boldsymbol{Q}_{1}^{1},\boldsymbol{Q}_{1}^{2}=\text{Split}(\boldsymbol{Q}_{1}),\boldsymbol{K}_{1}^{1},\boldsymbol{K}_{1}^{2}=\text{Split}(\boldsymbol{K}_{1}),\boldsymbol{V}_{1}^{1},\boldsymbol{V}_{1}^{2}=\text{Split}(\boldsymbol{V}_{1}),\\ &\boldsymbol{Q}_{2}^{1},\boldsymbol{Q}_{2}^{2}=\text{Split}(\boldsymbol{Q}_{2}),\boldsymbol{K}_{2}^{1},\boldsymbol{K}_{2}^{2}=\text{Split}(\boldsymbol{K}_{2}),\boldsymbol{V}_{2}^{1},\boldsymbol{V}_{2}^{2}=\text{Split}(\boldsymbol{V}_{2}).\\ \end{split}$ (5)
Then, $\boldsymbol{Q}_{1}^{1}$,$\boldsymbol{Q}_{2}^{1}$,$\boldsymbol{K}_{1}^{1}$,$\boldsymbol{K}_{2}^{1}$,$\boldsymbol{V}_{1}^{1}$,$\boldsymbol{V}_{2}^{1}$ are partitioned into horizontal windows and $\boldsymbol{Q}_{1}^{2}$,$\boldsymbol{Q}_{2}^{2}$,$\boldsymbol{K}_{1}^{2}$,$\boldsymbol{K}_{2}^{2}$,$\boldsymbol{V}_{1}^{2}$,$\boldsymbol{V}_{2}^{2}$ are partitioned into vertical windows, respectively. As shown in Fig. 7, given the size of the rectangular window is $[h,w]$, the input features are partitioned into non-overlapping rectangular patches, and the size of each patch is $[h,w,\frac{C}{2}]$. In the case of the horizontal partition, the width $w$ is greater than the height $h$ whereas in the case of the vertical partition the height $h$ is greater than the width $w$. Next, these windows are fed into the multi-head cross-attention (MCA) module to integrate the features between the RGB modality and HSI modality. More precisely, the cross-attention is calculated as
@@
Multi-head attention mechanism is also adopted in the RCA module, which indicates that we employ several groups of parameters to conduct the cross-attention, and the results are then combined to capture a more comprehensive and informative representation of the data.
@@
In summary, the input feature maps of HSI and RGB are divided into vertical and horizontal rectangular windows. Cross-attention is then applied between the vertical windows of the HSI and RGB feature maps, as well as between the horizontal windows. This rectangular attention mechanism enables the network to transfer the image details from the RGB modality to the HSI modality effectively and efficiently, leading to improved performance.
@@
Figure 8: Visualization results of synthesized HSIs which are generated under a single condition. Our proposed HSIGene supports image generation under single conditions, such as HED, segmentation, sketch, MLSD, content and text.
4 Experiments
4.1 Experimental Settings
4.1.1 Synthesis experiments setting
@@
The datasets used for HSI synthesis are listed in Table II and the details are illustrated in Sec. 3.1. Considering the spectral discrepancy between different datasets, we align their wavelength to 400-1000nm with 48 bands utilizing linear interpolation which could ensure spectral information preservation [48]. The datasets are cropped to size $256\times 256$ with stride 128, resulting in approximately 7k training images. For data augmentation, we employ the method introduced in Sec. 3 to obtain high-resolution images of the Heihe dataset with upscale factor ×$\times 2$ and Chikusei dataset with upscale factor ×$\times 4$. By cropping the super-resolved images of size $256\times 256$ with stride 128, we obtain 40k training images in total. To validate the robustness and generalization ability of our model, farmland images from the AID [49] dataset are used for evaluation, which exhibits scenes similar to real HSIs [19]. The AID dataset is also cropped with a stride of 128, resulting in 1k image patches of size $256\times 256$ in total. For each patch, we generate control conditions using the method introduced in Sec. 3.1 which are subsequently employed to synthesize HSIs. Finally, the quality of the generated images is evaluated to assess the model’s ability to generate realistic HSI data.
@@
We use the AdamW optimizer [50] with a learning rate of $10^{-5}$ to train the HSI synthesis diffusion model for 100k iterations. The batch size is set as 16. Our model contains 1.5 billion parameters in total, ensuring desirable generation under scaling laws. In addition, we initialize the parameters of the UNet component from the pretrained weights [24], ensuring that the diffusion model adapts effectively and efficiently to HSI synthesis.
4.1.2 Super-resolution experiments setting
@@
Approximately 55k high-resolution remote sensing images obtained from Google Earth Engine [43] are used to train the DSRNet described in Sec.3.3.2. We use real HSI data (i.e., the 7k cropped image patches) and select the RGB bands as auxiliary guidance to train the RGAN model proposed in 3.3.3. The training pairs are obtained by downsampling existing real HSIs with scales of 2 and 4 using the strategy [51].
@@
We adopt the same training strategy of the HSI synthesis model for DSRNet. For training RGAN, the learning rate is set to $10^{-4}$ with AdamW optimizer and the learning rate finally decreases to $10^{-5}$ with a cosine annealing strategy [52]. The batch size is set as 2 and the total epoch number is set as 30. L1 loss is used to train the network. To evaluate the image quality of the augmented data, we randomly select 100 real images from the Heihe dataset and compare the quality of these images super-resolved by different methods. Six super-resolution methods including MCNet [26], SSPSR [28], Bi-3DQRNN [27], SwinIR [53], ESSAformer [29], DSTrans [54] are taken for comparison. For a fair comparison, all competitive models are trained in the same setting.
@@ 4.1.3 Evaluation metrics
4.2 HSI synthesis results
4.2.1 Qualitative results
@@
In this section, we demonstrate the effectiveness of our model in achieving both single-condition and multi-condition control. The results of single-condition generation are illustrated in Fig. 8. As shown in the figure, the generated images are consistent with the specified condition but also exhibit pixel-wise differences. For example, given the sketch of farmland, the generated images demonstrate highly similar structures while the image contents include buildings and wasteland. By employing conditions that indicate textural attributes, such as HED or sketch, the generated images successfully maintain the structure of the input conditions. On the other hand, when the control condition is related to the image content, the model generates images that align closely with the reference content and exhibit various structures. However, single-condition control fails to guarantee simultaneous control over both the structure and content of the generated images. In addition, the model under single-condition constraints is unable to integrate multiple forms of guidance to achieve more precise and coherent control.
@@
The results of multi-condition generation are illustrated in Fig. 9. Compared with single-condition generation, more conditions enable more sophisticated and accurate generation. For example, as shown in Fig. 9, given conditions of HED and content indicating farmland, the generated images not only align with the structural prerequisites defined by the HED condition but also accurately reflect the content requirement. The multi-condition generation capability of our model offers significant advantages in producing images that are both structurally accurate and content-specific. Our model effectively handles different modalities and generates high-quality images, verifying its effectiveness and robustness to various conditions.
@@
Additionally, the visualization comparison between the generated and real spectral profiles is provided in Fig. 10. We select some generated HSIs whose content is similar to real HSIs, and for these images, we plot the spectral profiles at the same point for analysis. It can be seen that the generated spectral profiles are consistent with real spectral profiles, proving the effectiveness and usability of our proposed model in HSI synthesis. We also provide t-SNE [66] visualizations of the spectra generated by UBF [19], UnmixDiff [20], and our proposed method as shown in Fig. 11. It can be seen that the distribution of spectra generated by our method covers the real distribution more effectively. UBF generates spectra with significant differences from the real spectra due to the discrepancy between RGB modality image content and HSI content. UnmixDiff, on the other hand, generates abundance maps rather than directly generating HSIs, leading to performance limitations imposed by the unmixing network. Furthermore, UnmixDiff has a smaller model size compared to our model, which restricts its generation capabilities. In contrast, the latent diffusion technique that we used enables us to train a larger model, thus leading to better performance.